OpenStudy (anonymous):

Which is a solution to the equation 5x – y = –3? a.2,-7 b.1,2 c.-2,7 d.-2,-7

undeadknight26 (undeadknight26):

@helpme1.2 @jdoe0001 guys my brains gone out. I learned this but forgot it...

OpenStudy (jdoe0001):

well.... you have choices of ordered pairs, that is (x, y) say the 1st one is (2, -7) x = 2, y = -7 so let's try that one, let's plug it in our equation 5 x - y = -3 5(2) - (-7) = -3 10+7 \(\ne\) - 3 so is not a) now try the others

OpenStudy (jdoe0001):

when the left side matches the right side, then that's the solution

undeadknight26 (undeadknight26):

Oh i think i learned it by doing something like this: 5x - (0) = -3 and then 5(0) - y = -3

undeadknight26 (undeadknight26):

school confuses u...

OpenStudy (anonymous):

homeschooling confuses me

OpenStudy (anonymous):

you could also solve for each variable individually. isolate each variable one at a time and solve :) but @jdoe0001's way works too :) since you have answer choices, just sub in :)

OpenStudy (the_fizicx99):

^^ Just plug in the values and it should give you -3

OpenStudy (anonymous):

ok thx :$

OpenStudy (the_fizicx99):

Why the ":S"